OpenStudy (hhopke):

Circles. WILL MEDAL AND FAN!!! At what point in the first quadrant does the line with equation y = 0.5 x + 2 intersect the circle with radius 3 and center (0, 2)?

OpenStudy (hhopke):

The homework is due in 6 minutes, by the way, so I need the crash course version. :)

OpenStudy (hhopke):

@Directrix @jim_thompson5910 @Nnesha ? You guys are all really good at math and online.

jimthompson5910 (jim_thompson5910):

`circle with radius 3 and center (0, 2)` means we have the equation (x-h)^2 + (y-k)^2 = r^2 (x-0)^2 + (y-2)^2 = 3^2 x^2 + (y-2)^2 = 9

OpenStudy (hhopke):

right.

jimthompson5910 (jim_thompson5910):

replace the y with 0.5x+2, since y = 0.5x+2 x^2 + (y-2)^2 = 9 x^2 + (0.5x+2-2)^2 = 9 x^2 + (0.5x)^2 = 9 x^2 + 0.25x^2 = 9 I'll let you solve for x. Tell me what you get

OpenStudy (hhopke):

oooohhh...

OpenStudy (hhopke):

2.6832816

jimthompson5910 (jim_thompson5910):

plus or minus that, but yes

OpenStudy (hhopke):

thanks!

jimthompson5910 (jim_thompson5910):

then you use that x value to find the corresponding y value

jimthompson5910 (jim_thompson5910):

oh wait, you're in the first quadrant so nvm the negative part

OpenStudy (hhopke):

right. Thank you! :) (did it on a graphing calculator so I got the x and y value at the same time.)

jimthompson5910 (jim_thompson5910):

ok great, you're welcome